LHC 2011 - latest news

Week 27: Machine coordinators: Mike Lamont, Barbara Holzer (Saturday, Sunday only)

Major power cut 11:30 Sunday 10th July. Serious knock-on effects across accelerator complex. LHC cryogenics should be back late Monday/early Tuesday. Atlas and CMS solenoids not back until Thursday at earliest.

Planning - after recovery

  1. Verification tests:
    1. TCDQ and TCSG relative position --> re-measurement required.
    2. Beam presence flag intensity limit
    3. Beam Dump (inject and dump mode)
    4. Asynchronous dump 450 GeV B2 (pilot or 1-2 nominal bunches)
  2. Pilot + 2 nominal bunches (ATLAS needs the pilot on pilot collisions --> no overinjection)
    1. loss maps at injection
    2. 1 hour in stable beams
    3. loss maps at 3.5 TeV
  3. 48 bunches per beam
    •  new fill pattern for satellite on main collisions in ALICE: pilot over-injected, 250ns SPS gaps
    • MKI pulse length might need increasing
    • keep in stable beams for 3-4 hours (check with experiments)
  4. Default is to continue with the satellite collision scheme (unless requested otherwise by ALICE):
    1. 50ns_264b_249_0_240_36bpi8inj (1 fill)
    2. 50ns_840b_807_0_816_108bpi12inj (1 fill)
    3. 50ns_1380b_1331_0_1320_144bpi12inj.txt (luminosity production)

LHC coordination vistar



Link to report from morning meeting

Monday 11th July
  • Recovery still on-going
Sunday, 10th July
  • 11:30 Major power cut with knock-on effects across the CERN complex
  • Cryogenics estimated to be back late Monday/early Tuesday
09:00 meeting
Saturday, 9th July
  • 08:00 cryo back & stable, machine cold, QPS problem in 81, expert studying the problem


09:00 meeting

Monday, 4th July

  • 06:00 Start of Technical Stop

08:30 meeting

MD report